# Resonant Type-II + Type-I Hybrid Leptogenesis in SO(10) with Triplet–Neutrino Mass Degeneracy $M_{T}\simeq M_{N_{3}}$

### v1

@@ VI Conclusions
The resonant hybrid regime identified here is theoretically well motivated by the structure of SO(10) unification, phenomenologically rich, and highly predictive. It therefore provides a compelling and falsifiable pathway to baryogenesis, linking the origin of the baryon asymmetry to neutrino masses and low-energy CP violation within a unified framework. Finally, we have shown that although the heavy states responsible for hybrid leptogenesis reside far above the kinematic reach of the LHC, their effects do not completely decouple from collider observables. The hybrid CP-violating invariant induces a characteristic pattern of dimension-six operators, leading to correlated signatures in Higgs CP observables, Higgs–lepton couplings, lepton-flavour-violating dilepton production, and electroweak precision measurements. As illustrated in Fig. 3, these observables exhibit different decoupling behaviours with the heavy scale, highlighting the complementarity of Higgs, flavour, and precision probes in testing the hybrid leptogenesis framework. In particular, CP-odd observables in $h\to\tau^{+}\tau^{-}$ provide the most sensitive indirect collider window on the hybrid CP phase, establishing a direct conceptual link between the origin of the baryon asymmetry and precision measurements at the HL-LHC.

### v2

@@ V Numerical Results
Figure 5: Hybrid resonant leptogenesis in the plane of the hybrid CP phase $\phi_{\rm HR}$ and the heavy scale $\Lambda\sim M_{T}\sim M_{N_{3}}$. Left panel: colour shading shows a proxy for the baryon asymmetry $Y_{B}$ generated through the resonant interference of type-I and type-II contributions; the horizontal band indicates the region compatible with the observed baryon asymmetry. Right panel: indirect collider sensitivity to the same hybrid CP phase through CP-odd Higgs observables, illustrating the link between high-scale baryogenesis and precision measurements at the LHC.
Figure 5 provides a unified view of the hybrid leptogenesis mechanism and its experimental implications. The left panel displays the generation of the baryon asymmetry in the $\phi_{\rm HR}$$\Lambda$ plane, showing that successful leptogenesis occurs naturally for quasi-degenerate triplet and right-handed neutrino masses in the presence of a non-vanishing hybrid CP phase. The horizontal band highlights the region compatible with the observed baryon asymmetry, demonstrating that the hybrid resonant mechanism efficiently reproduces the correct order of magnitude without extreme parameter tuning.
The right panel shows the indirect collider sensitivity to the same hybrid CP phase through CP-odd Higgs observables induced by dimension-six operators after integrating out the heavy states. Although the leptogenesis scale lies far above the collider energy, the figure illustrates that sizable CP-violating effects persist over a wide range of the parameter space and can be probed through precision Higgs measurements at the LHC and its high-luminosity upgrade. Taken together, the two panels establish a direct conceptual link between the origin of the baryon asymmetry and experimentally accessible CP-violating observables.
Figure 6: Hybrid resonant leptogenesis parameter space in the plane of the hybrid CP phase $\phi_{\rm HR}$ and the degeneracy parameter $|\delta|=|M_{T}-M_{N_{3}}|/M_{N_{3}}$. The red island indicates the region where the resonant interference between type-I and type-II contributions generates a baryon asymmetry of the observed order. The hatched region is excluded by electric dipole moment constraints, while the blue contour denotes the approximate sensitivity of indirect collider CP-odd observables. The surviving overlap highlights the simultaneous viability and testability of the hybrid leptogenesis framework.
Figure 6 summarizes the phenomenological implications of the hybrid leptogenesis scenario in a single parameter-space scan. The finite red island arises from the combined requirements of resonant enhancement, quasi-degeneracy of the heavy states, and a non-vanishing hybrid CP phase, leading to a predictive region compatible with the observed baryon asymmetry. Electric dipole moment constraints exclude part of the parameter space but do not eliminate the viable region, while indirect collider sensitivity through CP-odd Higgs observables overlaps significantly with the baryogenesis-favoured domain. This demonstrates that hybrid resonant leptogenesis remains both viable and experimentally testable despite originating at very high scales.
Figure 7: Scaling of the baryon asymmetry with the degeneracy parameter $|\delta|$. Dots represent numerical scan points, while the solid line shows a power-law fit, illustrating the resonant enhancement as the heavy states approach degeneracy.
Figure 8: Linear correlation between the CP-odd Higgs observable and $\sin\phi_{\rm HR}$. The solid line represents a linear fit to the numerical points, demonstrating that the same hybrid CP phase governing leptogenesis induces approximately linear effects in collider observables.
Figure 9: Renormalization-group evolution of the effective CP-violating operator coefficient. Dots denote the evolved values at different scales, while the solid line shows a logarithmic fit, indicating mild RG enhancement of low-energy CP-odd observables.
Figure 7 shows the dependence of the baryon asymmetry on the degeneracy parameter $|\delta|=|M_{T}-M_{N_{3}}|/M_{N_{3}}$ obtained from a numerical scan. The discrete points correspond to individual parameter choices, while the solid line represents a power-law fit. The approximate scaling behaviour reflects the resonant enhancement of the CP asymmetry as the heavy states become increasingly degenerate, providing a quantitative characterization of the resonance structure underlying hybrid leptogenesis. Figure 8 illustrates the correlation between the CP-odd Higgs observable and $\sin\phi_{\rm HR}$. The numerical points are well described by a linear fit, indicating that the dominant contribution to the observable arises from the hybrid CP-violating phase. This behaviour supports the interpretation that the same CP phase responsible for baryogenesis induces approximately linear CP-violating effects in Higgs processes accessible at the LHC. Figure 9 displays the renormalization-group evolution of the effective CP-violating operator coefficient between the leptogenesis scale and the electroweak scale. The fitted logarithmic curve captures the leading RG behaviour and shows that the running induces only a mild enhancement of the coefficient. This justifies the use of leading-logarithmic RG evolution in estimating low-energy CP-odd observables while preserving their dependence on the hybrid CP phase.
@@ VI Conclusions
The resonant hybrid regime identified here is theoretically well motivated by the structure of SO(10) unification, phenomenologically rich, and highly predictive. It therefore provides a compelling and falsifiable pathway to baryogenesis, linking the origin of the baryon asymmetry to neutrino masses and low-energy CP violation within a unified framework. Finally, we have shown that although the heavy states responsible for hybrid leptogenesis reside far above the kinematic reach of the LHC, their effects do not completely decouple from collider observables. The hybrid CP-violating invariant induces a characteristic pattern of dimension-six operators, leading to correlated signatures in Higgs CP observables, Higgs–lepton couplings, lepton-flavour-violating dilepton production, and electroweak precision measurements. As illustrated in Fig. 3, these observables exhibit different decoupling behaviours with the heavy scale, highlighting the complementarity of Higgs, flavour, and precision probes in testing the hybrid leptogenesis framework. Figure 5 highlights how the same hybrid CP-violating invariant that generates the baryon asymmetry also controls experimentally accessible CP-odd Higgs observables, thereby establishing a direct link between high-scale leptogenesis and precision collider tests. In particular, CP-odd observables in $h\to\tau^{+}\tau^{-}$ provide the most sensitive indirect collider window on the hybrid CP phase, establishing a direct conceptual link between the origin of the baryon asymmetry and precision measurements at the HL-LHC.
The shaped parameter-space island in Fig. 6 shows that the hybrid CP phase responsible for the baryon asymmetry simultaneously governs EDM constraints and collider-accessible CP-odd observables, rendering the mechanism predictive and testable. The fitted scaling of the baryon asymmetry with the degeneracy parameter quantifies the resonant enhancement underlying the hybrid leptogenesis mechanism.
The approximately linear correlation between CP-odd Higgs observables and the hybrid CP phase demonstrates that collider measurements provide a direct probe of the CP violation responsible for baryogenesis.
The mild logarithmic renormalization-group evolution of the effective CP-violating operators shows that the qualitative link between high-scale leptogenesis and low-energy observables remains robust under running effects.